# Finite Element Operator Network for Solving Parametric PDEs

### v2

@@ Abstract
Partial differential equations (PDEs) underlie our understanding and prediction of natural phenomena across numerous fields, including physics, engineering, and finance. However, solving parametric PDEs is a complex task that necessitates efficient numerical methods. In this paper, we propose a novel approach for solving parametric PDEs using a Finite Element Operator Network (FEONet). Our proposed method leverages the power of deep learning in conjunction with traditional numerical methods, specifically the finite element method, to solve parametric PDEs in the absence of any paired input-output training data. We performed various experiments on several benchmark problems and confirmed that our approach has demonstrated excellent performance across various settings and environments, proving its versatility in terms of accuracy, generalization, and computational flexibility. Our FEONet framework shows potential for application in various fields where PDEs play a crucial role in modeling complex domains with diverse boundary conditions and singular behavior. Furthermore, we provide theoretical convergence analysis to support our approach, utilizing finite element approximation in numerical analysis.
keywords:
scientific machine learning, finite element methods, physics-informed operator learning, parametric PDEs, boundary layer
{MSCcodes}
65M60, 65N30, 68T20, 68U07
@@ 1 Introduction
In this regard, operator networks, an emerging field in research, address challenges by employing data-driven approaches to comprehend mathematical operators that govern physical systems. This is particularly relevant in solving parametric PDEs, as referenced in multiple studies [34, 29, 49, 7, 31, 35, 39, 16, 27, 41]. In particular, based upon the Universal Approximation Theorem for operators, [34] introduced the Deep Operator Network (DeepONet) architecture. These methods enable us to make fast predictions of solutions in real-time whenever the given PDE data changes, offering a new framework for solving parametric PDEs. However, these methods rely on training data for supervised learning that consists of pre-computed (either analytically or numerically) pairs of PDE parameters (e.g., external force, boundary condition, coefficients, initial condition) and their corresponding solutions. In general, generating a reliable training dataset requires extensive numerical computations, which can be computationally inefficient and time-consuming. Obtaining a sufficiently large dataset is also challenging, especially for systems with complex geometries or nonlinear equations. To overcome these limitations, the integration of PINNs and operator learning has led to the development of new methods, such as the Physics-Informed Neural Operator (PINO) [30] and Physics-Informed DeepONet (PIDeepONet) [49]. These methods strive to leverage the strengths of both PINNs and operator learning by incorporating physical equations into the neural operator’s loss function. However, they exhibit certain drawbacks, such as relatively low accuracy in handling complex geometries and difficulties in effectively addressing stiff problems like boundary layer issues, which are critical in real-world applications. Moreover, PIDeepONet tends to exhibit high generalization errors when lacking sufficient input samples. Additionally, employing neural networks as the solution space presents challenges in imposing boundary values, potentially leading to less accurate solutions.
@@
${u_{h}}({\bm{x}})=\sum\alpha_{k}\phi_{k}({{\bm{x}}}),\quad\bm{x}\in\mathbb{R}^{d},$ (1)
where ${u_{h}}$ is the FEM solution for the given PDEs. Motivated by (1), the FEONet can predict numerical solutions of the PDEs when given the initial conditions, external forcing functions, PDE coefficients, or boundary conditions as inputs. Therefore, the FEONet can learn multiple instances of the solutions of PDEs, making it a versatile approach for solving different types of PDEs in complex domains; see e.g. Figure 1. The loss function of the FEONet is designed based on the residual quantity of the finite element approximation, motivated by the classical FEM [11, 8]. This allows the FEONet to accurately approximate the solutions of PDEs, while also ensuring that the boundary conditions are exactly satisfied. This approach involves inferring coefficients, denoted as $\widehat{\alpha}_{k}$, which are subsequently utilized to construct the linear combination $\sum\widehat{\alpha}_{k}\phi_{k}$ to approximate the solution of PDEs. Inheriting the capability of the FEM handling boundary conditions, the predicted solution will also satisfy the exact boundary condition. Additionally, since we have adopted the FEM as a baseline framework, we can expect that our proposed method is robust, reliable and accurate. It is noteworthy that due to the intrinsic structure of the proposed FEONet scheme, we do not need any paired input-output training data and hence the model can be trained in an unsupervised manner.
@@ 2 Finite element operator network (FEONet)
$-\varepsilon\,{\rm{div}}\,(\bm{a}(\bm{x})\nabla u)+\mathcal{F}(u)=f\quad{\rm{in}}\,\,D.$ (2)
@@ 2.1 Model description
We begin with the following variational formulation of the given PDE (2): find $u\in V$ satisfying
$B[u,v]:=\varepsilon\int_{D}\bm{a}(\bm{x})\nabla u\cdot\nabla v\,\mathrm{d}\bm{x}+\int_{D}\mathcal{F}(u)v\,\mathrm{d}\bm{x}=\int_{D}fv\,\mathrm{d}\bm{x}=:\ell(v)\quad{\text{for all}}\,\,v\in V,$ (3)
@@
$u_{h}=\sum^{N(h)}_{k=1}\alpha_{k}\phi_{k},\quad\alpha_{i}\in\mathbb{R},$ (5)
the Galerkin approximation (4) is transformed into the linear algebraic system
$A\alpha=F\quad{\text{with}}\,\,A_{ik}:=B[\phi_{k},\phi_{i}]\,\,{\rm{and}}\,\,F_{i}:=\ell(\phi_{i}),$ (6)
where the matrix $A\in\mathbb{R}^{N(h)\times N(h)}$ is invertible provided that the given PDEs have a suitable structure. Therefore, we can determine the coefficients $\{\alpha_{k}\}^{N(h)}_{k=1}$ by solving the system of linear equations (6), and consequently, produce the approximate solution $u_{h}\in V_{h}$ as defined in (5).
@@
Now, in this setting let us describe the method we propose in this paper. In our scheme, an input for the neural network is the data for the given PDEs. In this paper, we choose external force as a prototype input feature. Note, however, that the same scheme can also be developed with other types of data, e.g., boundary conditions, diffusion coefficients or initial conditions for time-dependent problems. See, for example, Section 3.2 where extensive experiments were performed addressing various types of input data. For this purpose, the external forcing terms are parametrized by the random parameter $\bm{\omega}$ contained in the (possibly high-dimensional) parameter space $\Omega$. For each realization $f(\bm{x};\bm{\omega})$ (and hence for each load vector $F(\bm{\omega})$ defined in (6)), instead of computing the coefficients from the linear algebraic system (6), we approximate the coefficients $\alpha$ by a deep neural network. More precisely, these input features representing external forces pass through the deep neural network and the coefficients $\{\widehat{\alpha}_{k}\}$ are generated as an output of the neural network. We then reconstruct the solution
$\widehat{u}_{h}(\bm{x};\bm{\omega})=\sum_{k=1}^{N(h)}\widehat{\alpha}_{k}(\bm{\omega})\phi_{k}(\bm{x}).$ (7)
@@
For the training, we use a residual of the variational formulation (4) and define the loss function by summing over all basis functions: for randomly chosen parameters $\bm{\omega}_{1},\cdots,\bm{\omega}_{M}\in\Omega$,
$\mathcal{L}^{M}(\widehat{u}_{h})=\frac{1}{M}\sum^{M}_{m=1}\sum^{N(h)}_{i=1}|B[\widehat{u}_{h}(\bm{x};\bm{\omega}_{m}),\phi_{i}(\bm{x})]-\ell(\phi_{i}(\bm{x});\bm{\omega}_{m})|^{2}.$ (8)
@@
For each training epoch, once the neural network parameters are updated in the direction of minimizing the loss function (8), the external force $f(\bm{x};\bm{\omega})$ passes through this updated neural network to generate more refined coefficients, and this procedure is repeated until the sufficiently small loss is achieved. A schematic diagram of the FEONet algorithm is depicted in Figure 2.
@@ 2.2 Convergence analysis of FEONet
In this section, we address the convergence result of the FEONet which provides theoretical justification for the proposed method. As outlined in the previous subsection, the proposed method is based on the FEM, and the finite element approximation $u_{h}$ in (5) can be considered as an intermediate solution between the true solution $u$ of (3) and the approximate solution $\widehat{u}_{h}$ predicted by the FEONet.
@@
where the coefficient $\bm{a}(\bm{x})$ is uniformly elliptic and $c(\bm{x})\geq 0$. In the present paper, we focus on the case when the finite element parameter $h>0$ is fixed and deal with the convergence of the FEONet prediction to the finite element solution $u_{h}$. The analysis for general equations that cannot be covered by the form (9) and the theoretical analysis addressing the role of $h>0$ in the error estimate are of independent interest and will be investigated in the forthcoming paper.
@@
$f(\bm{x};\bm{\omega})\in C(\Omega;L^{1}(D))\quad{\rm{and}}\quad\sup_{\bm{\omega}\in\Omega}\int_{D}|f(\bm{x};\bm{\omega})|\,\mathrm{d}\bm{x}<C$ (10)
@@
$B[u,v]:={\int_{D}[\bm{a}(\bm{x})\nabla u\cdot\nabla v+c(\bm{x})uv]\,\mathrm{d}\bm{x}}=\int_{D}f(\bm{x})v\,\mathrm{d}\bm{x}=:\ell(v)\quad\forall v\in H^{1}_{0}(D).$ (11)
@@
$u_{h}(\bm{x},\bm{\omega})=\sum^{N(h)}_{k=1}\alpha^{*}_{k}(\bm{\omega})\phi_{k}(\bm{x}),$ (13)
@@
$A_{ik}=B[\phi_{k},\phi_{i}]\quad{\rm{and}}\quad F_{i}=\ell(\phi_{i}).$ (15)
In other words, $u_{h}$ in (13) is the finite element approximation of the solution to the equation (9), as described in Section 2.1. It is noteworthy that instead of the equations (14), $\alpha^{*}$ can also be characterized in the following way:
$\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in C(\Omega,\mathbb{R}^{N(h)})}\mathcal{L}(\alpha),$ (16)
@@
$\mathcal{L}(\alpha)=\mathbb{E}_{\bm{\omega}\sim\mathbb{P}_{\Omega}}\bigg{[}\sum^{N(h)}_{i=1}|B[\widehat{u}(\bm{\omega}),\phi_{i}]-\ell(\phi_{i};(\bm{\omega}))|^{2}\bigg{]}=\|A\alpha(\bm{\omega})-F(\bm{\omega})\|^{2}_{L^{2}(\Omega)}.$ (17)
@@
$\Theta_{\vec{\bm{n}}}=\left\{\{(W^{\ell},b^{\ell})\}^{L}_{\ell=1}:W^{\ell}\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}},\,b^{\ell}\in\mathbb{R}^{n_{\ell}}\right\}.$ (19)
@@
$\mathcal{N}^{*}_{n}=\{\mathcal{R}[\theta]:\theta\in\Theta_{\vec{\bm{n}}_{n}}\}.$ (20)
@@
$\lim_{n\rightarrow\infty}\inf_{\hat{g}\in\mathcal{N}^{*}_{n}}\|\hat{g}-g\|_{C(K)}=0.$ (21)
In the present paper, the target function to be approximated by neural networks is $\alpha^{*}(\bm{\omega})$. Due to the assumption (10), we can verify that $\alpha^{*}(\bm{\omega})$ is continuous and bounded. This motivates us to deal with the family of bounded neural networks as an ansatz space. For this, let us write $\sup_{\bm{\omega}\in\Omega}|\alpha^{*}(\bm{\omega})|<C_{\alpha}$, and assume that $\sigma^{*}$ is a bounded activation function with continuous inverse (e.g., sigmoid or tanh activation functions) with $\sigma_{-}:=\inf_{x\in\mathbb{R}}\sigma^{*}(x)$ and $\sigma_{+}:=\sup_{x\in\mathbb{R}}\sigma^{*}(x)$. We then define a function $h:[\sigma_{-},\sigma_{+}]\rightarrow[-C_{\alpha},C_{\alpha}]$ satisfying $h(\sigma_{-})=-C_{\alpha}$, $h(\sigma_{+})=C_{\alpha}$. By the assumptions, $h$ also has a continuous inverse. With the above notations, we consider the following family of neural networks:
$\mathcal{N}_{n}=\left\{h\circ\sigma^{*}(g_{n}):g_{n}\in\mathcal{N}_{n}^{*}\right\},$ (22)
@@
$\|g\|_{C(\Omega)}\leq C_{\alpha},\quad\forall g\in\bigcup_{n\in\mathbb{N}}\mathcal{N}_{n}.$ (23)
@@
Let $\Omega$ be a compact set and $\alpha^{*}:\Omega\rightarrow\mathbb{R}^{N-1}$ is a continuous function defined in (14). Then we have
$\lim_{n\rightarrow\infty}\inf_{\tilde{g}\in\mathcal{N}_{n}}\|\tilde{g}-\alpha^{*}\|_{C(\Omega)}=0.$ (24)
@@
$\widehat{\alpha}(n)=\operatorname*{arg\,min}_{\alpha\in\mathcal{N}_{n}}\mathcal{L}(\alpha),$ (25)
@@
$u_{h,n}(\bm{x};\bm{\omega})=\sum^{N(h)}_{k=1}\widehat{\alpha}(n)_{k}(\bm{\omega})\phi_{k}(\bm{x}).$ (26)
@@
$\widehat{\alpha}(n,M)=\operatorname*{arg\,min}_{\alpha\in\mathcal{N}_{n}}\mathcal{L}^{M}(\alpha),$ (27)
@@
$\mathcal{L}^{M}(\alpha)=\frac{|\Omega|}{M}\sum_{m=1}^{M}\sum^{N(h)}_{i=1}|B[\widehat{u}(\bm{\omega}_{m}),\phi_{i}]-\ell(\phi_{i};(\bm{\omega}_{m}))|^{2}=\frac{|\Omega|}{M}\sum^{M}_{m=1}|A\alpha(\bm{\omega}_{m})-F(\bm{\omega}_{m})|^{2},$ (28)
@@
$u_{h,n,M}(\bm{x};\bm{\omega})=\sum^{N(h)}_{k=1}\widehat{\alpha}(n,M)_{k}(\bm{\omega})\phi_{k}(\bm{x}),$ (29)
which is the numerical solution actually computed by the scheme proposed in the present paper. In this paper, we assume that it is possible to find the exact minimizers for the problems (25) and (27), and the error that occurred by the optimization can be ignored.
@@
$\|u-u_{h,n,M}\|_{L^{2}(\Omega;L^{2}(D))}\rightarrow 0\quad{\rm{as}}\,\,h\rightarrow 0\,\,{\rm{and}}\,\,n,M\rightarrow\infty.$ (30)
@@
$u-u_{h,n,M}=(u-u_{h})+(u_{h}-u_{h,n})+(u_{h,n}-u_{h,n,M}).$ (31)
The first error is the one caused by the finite element approximation, which is assumed to be negligible for a proper choice of $h$. In fact, according to the classical theory of the FEM (e.g. [8, 11]) if $h>0$ is sufficiently small, then we can expect to obtain a sufficiently small error. Therefore, in this paper, we shall assume that we have chosen suitable $h>0$ which guarantees a sufficiently small error for the finite element approximation and only focus on the convergence of $u_{h}-u_{h,n,M}$ with a fixed $h>0$.

### v3

# Finite Element Operator Network for Solving Elliptic-type Parametric PDEs

@@ Abstract
Partial differential equations (PDEs) underlie our understanding and prediction of natural phenomena across numerous fields, including physics, engineering, and finance. However, solving parametric PDEs is a complex task that necessitates efficient numerical methods. In this paper, we propose a novel approach for solving parametric PDEs using a Finite Element Operator Network (FEONet). Our proposed method leverages the power of deep learning in conjunction with traditional numerical methods, specifically the finite element method, to solve parametric PDEs in the absence of any paired input-output training data. We performed various experiments on several benchmark problems and confirmed that our approach has demonstrated excellent performance across various settings and environments, proving its versatility in terms of accuracy, generalization, and computational flexibility. While our method is not meshless, the FEONet framework shows potential for application in various fields where PDEs play a crucial role in modeling complex domains with diverse boundary conditions and singular behavior. Furthermore, we provide theoretical convergence analysis to support our approach, utilizing finite element approximation in numerical analysis.
Keywords: scientific machine learning, finite element methods, physics-informed operator learning, parametric PDEs, boundary layer
AMS Classification: 65M60, 65N30, 68T20, 68U07
@@ 1 Introduction
In this regard, operator networks, an emerging field in research, address challenges by employing data-driven approaches to comprehend mathematical operators that govern physical systems. This is particularly relevant in solving parametric PDEs, as referenced in multiple studies [37, 32, 53, 7, 34, 38, 42, 16, 30, 44]. In particular, based upon the Universal Approximation Theorem for operators, [37] introduced the Deep Operator Network (DeepONet) architecture. Moreover, the authors in [38] proposed an extension of the DeepONet model, called POD-DeepONet, which replaces a part of the DeepONet architecture with a prefixed basis obtained through proper orthogonal decomposition (POD) on the training data. These methods enable us to make fast predictions of solutions in real-time whenever the given PDE data changes, offering a new framework for solving parametric PDEs. However, these methods rely on training data for supervised learning that consists of pre-computed (either analytically or numerically) pairs of PDE parameters (e.g., external force, boundary condition, coefficients, initial condition) and their corresponding solutions. In general, generating a reliable training dataset requires extensive numerical computations, which can be computationally inefficient and time-consuming. Obtaining a sufficiently large dataset is also challenging, especially for systems with complex geometries or nonlinear equations. To overcome these limitations, the integration of PINNs and operator learning has led to the development of new methods, such as the Physics-Informed Neural Operator (PINO) [33] and Physics-Informed DeepONet (PIDeepONet) [53]. These methods strive to leverage the strengths of both PINNs and operator learning by incorporating physical equations into the neural operator’s loss function. However, they exhibit certain drawbacks, such as relatively low accuracy in handling complex geometries and difficulties in effectively addressing stiff problems like boundary layer issues, which are critical in real-world applications. Moreover, PIDeepONet tends to exhibit high generalization errors when lacking sufficient input samples. Additionally, employing neural networks as the solution space presents challenges in imposing boundary values, potentially leading to less accurate solutions.
@@
${u_{h}}({\bm{x}})=\sum\alpha_{k}\phi_{k}({{\bm{x}}}),\quad\bm{x}\in\mathbb{R}^{d},$ (1.1)
where ${u_{h}}$ is the FEM solution for the given PDEs. Motivated by (1.1), the FEONet can predict numerical solutions of the PDEs when given the initial conditions, external forcing functions, PDE coefficients, or boundary conditions as inputs. Therefore, the FEONet can learn multiple instances of the solutions of PDEs, making it a versatile approach for solving different types of PDEs in complex domains; see e.g. Figure 1. The loss function of the FEONet is designed based on the residual quantity of the finite element approximation, motivated by the classical FEM [11, 8]. This allows the FEONet to accurately approximate the solutions of PDEs, while also ensuring that the boundary conditions are exactly satisfied. This approach involves inferring coefficients, denoted as $\widehat{\alpha}_{k}$, which are subsequently utilized to construct the linear combination $\sum\widehat{\alpha}_{k}\phi_{k}$ to approximate the solution of PDEs. Inheriting the capability of the FEM handling boundary conditions, the predicted solution will also satisfy the exact boundary condition. Note further that since we reconstruct the solution as a linear combination of the coefficients predicted by neural networks and the basis function, the FEONet scheme generates a solution as a function defined on the whole domain, not only on particular points in the domain. Additionally, since we have adopted the FEM as a baseline framework, we can expect that our proposed method is robust, reliable, and accurate. It is noteworthy that due to the intrinsic structure of the proposed FEONet scheme, we do not need any paired input-output training data and hence the model can be trained in an unsupervised manner.
@@ 2 Finite element operator network (FEONet)
$-\varepsilon\,{\rm{div}}\,(\bm{a}(\bm{x})\nabla u)+\mathcal{F}(u)=f\quad{\rm{in}}\,\,D.$ (2.1)
@@ 2.1 Model description
We begin with the following variational formulation of the given PDE (2.1): find $u\in V$ satisfying
$B[u,v]:=\varepsilon\int_{D}\bm{a}(\bm{x})\nabla u\cdot\nabla v\,\mathrm{d}\bm{x}+\int_{D}\mathcal{F}(u)v\,\mathrm{d}\bm{x}=\int_{D}fv\,\mathrm{d}\bm{x}=:\ell(v)\quad{\text{for all}}\,\,v\in V,$ (2.2)
@@
$u_{h}=\sum^{N(h)}_{k=1}\alpha_{k}\phi_{k},\quad\alpha_{i}\in\mathbb{R},$ (2.4)
the Galerkin approximation (2.3) is transformed into the linear algebraic system
$A\alpha=F\quad{\text{with}}\,\,A_{ik}:=B[\phi_{k},\phi_{i}]\,\,{\rm{and}}\,\,F_{i}:=\ell(\phi_{i}),$ (2.5)
where the matrix $A\in\mathbb{R}^{N(h)\times N(h)}$ is invertible provided that the given PDEs have a suitable structure. Therefore, we can determine the coefficients $\{\alpha_{k}\}^{N(h)}_{k=1}$ by solving the system of linear equations (2.5), and consequently, produce the approximate solution $u_{h}\in V_{h}$ as defined in (2.4).
@@
Now, in this setting let us describe the method we propose in this paper. In our scheme, an input for the neural network is the data for the given PDEs. In this paper, we choose external force as a prototype input feature. Note, however, that the same scheme can also be developed with other types of data, e.g., boundary conditions, diffusion coefficients, or initial conditions for time-dependent problems. See, for example, Section 3.1.4 where extensive experiments were performed addressing various types of input data. For this purpose, the external forcing terms are parametrized by the random parameter $\bm{\omega}$ contained in the (possibly high-dimensional) parameter space $\Omega$. For each realization $f(\bm{x};\bm{\omega})$ (and hence for each load vector $F(\bm{\omega})$ defined in (2.5)), instead of computing the coefficients from the linear algebraic system (2.5), we approximate the coefficients $\alpha$ by a deep neural network. More precisely, these input features representing external forces pass through the deep neural network and the coefficients $\{\widehat{\alpha}_{k}\}$ are generated as an output of the neural network. We then reconstruct the solution
$\widehat{u}_{h}(\bm{x};\bm{\omega})=\sum_{k=1}^{N(h)}\widehat{\alpha}_{k}(\bm{\omega})\phi_{k}(\bm{x}).$ (2.6)
@@
For the training, we use a residual of the variational formulation (2.3) and define the loss function by summing over all basis functions: for randomly chosen parameters $\bm{\omega}_{1},\cdots,\bm{\omega}_{M}\in\Omega$,
$\mathcal{L}^{M}(\widehat{u}_{h})=\frac{1}{M}\sum^{M}_{m=1}\sum^{N(h)}_{i=1}|B[\widehat{u}_{h}(\bm{x};\bm{\omega}_{m}),\phi_{i}(\bm{x})]-\ell(\phi_{i}(\bm{x});\bm{\omega}_{m})|^{2}.$ (2.7)
Note that our proposed loss function is based on the variational form of the given problem, and the parameters are updated in the direction that minimizes the corresponding loss function (2.7) defined above. In this paper, we assume that it is possible to find the exact minimizers, allowing us to ignore any errors arising from the optimization process. It is worth noting that loss functions based on the variational form have also been considered in previous work related to PINNs [24, 25]. However, unlike our approach, those methods were designed for single-instance predictions and did not utilize the P1 basis elements that are integral to our operator network. Instead, they employed different types of test functions. While [19] also proposes an operator-learning method and uses the variational form of governing equations as a loss function along with input-output data pairs, there is a difference from our FEONet approach, which is completely free from the need for input-output data pairs. The detailed algorithm for training FEONet is provided in the pseudo-code below.
Algorithm 1 FEONet Training Procedure
1:Training samples $f(\bm{x};\bm{\omega}_{m})$ for $m=1,\ldots,M$;
2:The coefficients $\widehat{\alpha}(\bm{\omega}_{m})=[\widehat{\alpha}_{k}(\bm{\omega}_{m})]_{k=1,\ldots,N(h)}$ for $m=1,\ldots,M$. Therefore, the predicted solution is $\widehat{u}_{h}(\bm{x};\bm{\omega}_{m})=\sum_{k=1}^{N(h)}\widehat{\alpha}_{k}(\bm{\omega}_{m})\phi_{k}(\bm{x})$;
3:Pre-compute $A$ ($A_{ik}=B[\phi_{k},\phi_{i}]$) and $F(\bm{\omega}_{m})$ for $m=1,\ldots,M$ using basis functions $\{\phi_{i}\}_{i=1}^{N(h)}$ based on domain triangulation $\{\mathcal{T}_{h}\}$;
4:for $n=1,2,\ldots,N_{\text{epochs}}$ do
5:     $\mathcal{L}^{M}\leftarrow 0$;
6:     for $m=1,\ldots,M$ do
7:         Compute $\widehat{\alpha}(\bm{\omega}_{m})$ using FEONet with input $f(\bm{x};\bm{\omega}_{m})$;
8:         $\mathcal{L}^{M}\leftarrow\mathcal{L}^{M}+|A\widehat{\alpha}(\bm{\omega}_{m})-F(\bm{\omega}_{m})|^{2}$;
9:     end for
10:     Update FEONet parameters by minimizing $\mathcal{L}^{M}$;
11:end for
@@
For each training epoch, once the neural network parameters are updated in the direction of minimizing the loss function (2.7), the external force $f(\bm{x};\bm{\omega})$ passes through this updated neural network to generate more refined coefficients, and this procedure is repeated until the sufficiently small loss is achieved. A schematic diagram of the FEONet algorithm is depicted in Figure 2.
@@
Remark 2.2.
There are several reasons why we used the variational residual (2.7) as a loss function instead of other types of loss functions, e.g., the physics-informed loss function. Since our method is inherited from the classical FEM, in order to use the well-known techniques from numerical analysis, it is more “natural” to consider the loss function we chose. As a consequence, we are able to successfully obtain the theoretical convergence result from the viewpoint of the FEM, where the analysis mainly depends on the form of the loss function. To be more specific, in appendix A.1 and our subsequent paper [20], we prove the best approximation property of the FEONet, providing the theoretical evidence that we chose the right loss function. In summary, we have adopted the variational residual as a loss function from the theoretical perspective. On the other hand, there are also some advantages of using this weak formulation as a loss function. For example, we can reduce the computation error for the derivative like the variational physics-informed neural networks [24, 25].
@@ 2.2 Convergence analysis of FEONet
In this section, we address the convergence result of the FEONet which provides theoretical justification for the proposed method. As outlined in the previous subsection, the proposed method is based on the FEM, and the finite element approximation $u_{h}$ in (2.4) can be considered as an intermediate solution between the true solution $u$ of (2.2) and the approximate solution $\widehat{u}_{h}$ predicted by the FEONet.
@@
where the coefficient $\bm{a}(\bm{x})$ is uniformly elliptic and $c(\bm{x})\geq 0$. In the present paper, we focus on the case when the finite element parameter $h>0$ is fixed and deal with the convergence of the FEONet prediction to the finite element solution $u_{h}$. The analysis for general equations that cannot be covered by the form (2.8) and the theoretical analysis addressing the role of $h>0$ in the error estimate are of independent interest and will be investigated in the forthcoming paper.
@@
$f(\bm{x};\bm{\omega})\in C(\Omega;L^{1}(D))\quad{\rm{and}}\quad\sup_{\bm{\omega}\in\Omega}\int_{D}|f(\bm{x};\bm{\omega})|\,\mathrm{d}\bm{x}<C$ (2.9)
@@
$B[u,v]:={\int_{D}[\bm{a}(\bm{x})\nabla u\cdot\nabla v+c(\bm{x})uv]\,\mathrm{d}\bm{x}}=\int_{D}f(\bm{x})v\,\mathrm{d}\bm{x}=:\ell(v)\quad\forall v\in H^{1}_{0}(D).$ (2.10)
@@
$u_{h}(\bm{x},\bm{\omega})=\sum^{N(h)}_{k=1}\alpha^{*}_{k}(\bm{\omega})\phi_{k}(\bm{x}),$ (2.12)
@@
$A_{ik}=B[\phi_{k},\phi_{i}]\quad{\rm{and}}\quad F_{i}=\ell(\phi_{i}).$ (2.14)
In other words, $u_{h}$ in (2.12) is the finite element approximation of the solution to the equation (2.8), as described in Section 2.1. It is noteworthy that instead of the equations (2.13), $\alpha^{*}$ can also be characterized in the following way:
$\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in C(\Omega,\mathbb{R}^{N(h)})}\mathcal{L}(\alpha),$ (2.15)
@@
$\mathcal{L}(\alpha)=\mathbb{E}_{\bm{\omega}\sim\mathbb{P}_{\Omega}}\bigg{[}\sum^{N(h)}_{i=1}|B[\widehat{u}(\bm{\omega}),\phi_{i}]-\ell(\phi_{i};(\bm{\omega}))|^{2}\bigg{]}=\|A\alpha(\bm{\omega})-F(\bm{\omega})\|^{2}_{L^{2}(\Omega)}.$ (2.16)
@@
$\Theta_{\vec{\bm{n}}}=\left\{\{(W^{\ell},b^{\ell})\}^{L}_{\ell=1}:W^{\ell}\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}},\,b^{\ell}\in\mathbb{R}^{n_{\ell}}\right\}.$ (2.18)
@@
$\mathcal{N}^{*}_{n}=\{\mathcal{R}[\theta]:\theta\in\Theta_{\vec{\bm{n}}_{n}}\}.$ (2.19)
@@
$\lim_{n\rightarrow\infty}\inf_{\hat{g}\in\mathcal{N}^{*}_{n}}\|\hat{g}-g\|_{C(K)}=0.$ (2.20)
In the present paper, the target function to be approximated by neural networks is $\alpha^{*}(\bm{\omega})$. Due to the assumption (2.9), we can verify that $\alpha^{*}(\bm{\omega})$ is continuous and bounded. This motivates us to deal with the family of bounded neural networks as an ansatz space. For this, let us write $\sup_{\bm{\omega}\in\Omega}|\alpha^{*}(\bm{\omega})|<C_{\alpha}$, and assume that $\sigma^{*}$ is a bounded activation function with continuous inverse (e.g., sigmoid or tanh activation functions) with $\sigma_{-}:=\inf_{x\in\mathbb{R}}\sigma^{*}(x)$ and $\sigma_{+}:=\sup_{x\in\mathbb{R}}\sigma^{*}(x)$. We then define a function $h:[\sigma_{-},\sigma_{+}]\rightarrow[-C_{\alpha},C_{\alpha}]$ satisfying $h(\sigma_{-})=-C_{\alpha}$, $h(\sigma_{+})=C_{\alpha}$. By the assumptions, $h$ also has a continuous inverse. With the above notations, we consider the following family of neural networks:
$\mathcal{N}_{n}=\left\{h\circ\sigma^{*}(g_{n}):g_{n}\in\mathcal{N}_{n}^{*}\right\},$ (2.21)
@@
$\|g\|_{C(\Omega)}\leq C_{\alpha},\quad\forall g\in\bigcup_{n\in\mathbb{N}}\mathcal{N}_{n}.$ (2.22)
@@
Let $\Omega$ be a compact set and $\alpha^{*}:\Omega\rightarrow\mathbb{R}^{N-1}$ is a continuous function defined in (2.13). Then we have
$\lim_{n\rightarrow\infty}\inf_{\tilde{g}\in\mathcal{N}_{n}}\|\tilde{g}-\alpha^{*}\|_{C(\Omega)}=0.$ (2.23)
@@
$\widehat{\alpha}(n)=\operatorname*{arg\,min}_{\alpha\in\mathcal{N}_{n}}\mathcal{L}(\alpha),$ (2.24)
@@
$u_{h,n}(\bm{x};\bm{\omega})=\sum^{N(h)}_{k=1}\widehat{\alpha}(n)_{k}(\bm{\omega})\phi_{k}(\bm{x}).$ (2.25)
@@
$\widehat{\alpha}(n,M)=\operatorname*{arg\,min}_{\alpha\in\mathcal{N}_{n}}\mathcal{L}^{M}(\alpha),$ (2.26)
@@
$\mathcal{L}^{M}(\alpha)=\frac{|\Omega|}{M}\sum_{m=1}^{M}\sum^{N(h)}_{i=1}|B[\widehat{u}(\bm{\omega}_{m}),\phi_{i}]-\ell(\phi_{i};(\bm{\omega}_{m}))|^{2}=\frac{|\Omega|}{M}\sum^{M}_{m=1}|A\alpha(\bm{\omega}_{m})-F(\bm{\omega}_{m})|^{2},$ (2.27)
@@
$u_{h,n,M}(\bm{x};\bm{\omega})=\sum^{N(h)}_{k=1}\widehat{\alpha}(n,M)_{k}(\bm{\omega})\phi_{k}(\bm{x}),$ (2.28)
which is the numerical solution actually computed by the scheme proposed in the present paper.
@@
$\|u-u_{h,n,M}\|_{L^{2}(\Omega;L^{2}(D))}\rightarrow 0\quad{\rm{as}}\,\,h\rightarrow 0\,\,{\rm{and}}\,\,n,M\rightarrow\infty.$ (2.29)
@@
$u-u_{h,n,M}=(u-u_{h})+(u_{h}-u_{h,n})+(u_{h,n}-u_{h,n,M}).$ (2.30)
The first error is the one caused by the finite element approximation, which is assumed to be negligible for a proper choice of $h$. In fact, according to the classical theory of the FEM (e.g. [8, 11]) if $h>0$ is sufficiently small, then we can expect to obtain a sufficiently small error. Therefore, in this paper, we shall assume that we have chosen suitable $h>0$ which guarantees a sufficiently small error for the finite element approximation, and only focus on the convergence of $u_{h}-u_{h,n,M}$ with a fixed $h>0$.